OpenStudy (anonymous):

solve the equation 3/4a=12

OpenStudy (anonymous):

3/4a = 12 Dividing fractions means multiplying by the reciprocal. a = 12(4/3) a = 48/3 a = 16

OpenStudy (anonymous):

a = 12 x 4/3 a = 48/3 so, a = 16. :)

OpenStudy (theviper):

3/4a=12 then,3=4a*12 then,3/12=4a then,3*4/12=a then,12/12=a=1.

OpenStudy (theviper):

@Mimi_x3 Am I right here?

OpenStudy (mimi_x3):

Does not seem right. This part does not make sense: \[\frac{3}{12} =4a => 3*\frac{4}{12} =4a\]

OpenStudy (theviper):

How can it be 3*4/12=4 as u have taken 4 @ right hand side.

OpenStudy (theviper):

Plz solve it step by step:)

OpenStudy (mimi_x3):

An easier way is: \[\frac{3}{4a} =12 => 3= 4a*12 => 3=48a => 48a=3 \]

OpenStudy (theviper):

Then Calcmathlete & Hally are wrong:/

OpenStudy (theviper):

What is the ans. @Damnicansing.

OpenStudy (mimi_x3):

Most likely; since WolframAlpha agrees with me: http://www.wolframalpha.com/input/?i=3%2F4a%3D12

OpenStudy (theviper):

I am seeing:)

OpenStudy (maheshmeghwal9):

hey dude! http://www.wolframalpha.com/input/?i=3%2F4x%3D12

OpenStudy (maheshmeghwal9):

just take x as a

OpenStudy (mimi_x3):

same thing lol.

OpenStudy (maheshmeghwal9):

ya but above is confusion for him/her lol so what is going, if we give one chance again^_^